# Investigating FRB 20240114A with FAST: Morphological Classification and Drifting Rate Measurements in a Burst-Cluster Framework

### v1

@@ 3.2 Upward Drifting burst-clusters
In Zhou et al. (2022), FRB 20201124A exhibited 263 downward drifting bursts (including single-, double-, and multi-components bursts) alongside 3 upward drifting bursts, all of which were double-components bursts. In contrast, our classification of FRB 20240114A shows that upward drifting burst-clusters account for 23.82% (233 out of 978) of the total burst-clusters exhibiting drifting behavior. However, the drifting mode of single-component burst-clusters is significantly affected by the choice of DM, while double- and multiple-components burst-clusters are not subject to this influence. If we exclude single-component burst-clusters with upward drifting, the number of upward drifting double- and multiple-components burst-clusters still reaches 91, accounting for 10.89% of the total sample of 836 burst-clusters exhibiting drifting behavior (excluding single-component burst-clusters with upward drifting). Furthermore, if we consider only upward drifting bursts, specifically those upward drifting burst-clusters consisting solely of consecutive time intervals, this number drastically decreases to just 9, as illustrated in Figure 6. This proportion is comparable to that observed in Zhou et al. (2022). Upward drifting bursts, though rare, find theoretical grounding in magnetospheric trigger dynamics. Sparks triggered earlier may occur along field lines with a larger radius of curvature (Wang et al., 2020). Propagation effects such as interstellar scintillation can enhance the detection of upward drifting features through temporary intensity amplification (Rickett, 1990), although intrinsic emission mechanisms remain dominant.
3.3 No Drifting burst-clusters
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Representative examples of burst-cluster classifications. The first row shows No Drifting burst-clusters: (a) ND-1, a single-component burst-cluster with a stable emission band and no frequency drift, (b) ND-2, a double-components burst-cluster with two components showing no drift, and (c) ND-m, a multiple burst-cluster with parallel components. The second row illustrates burst-clusters with no evidence for drifting: (d) NE-H, a high-frequency burst-cluster with insufficient bandwidth to determine drift, (e) NE-L, a low-frequency burst-cluster with insufficient bandwidth to determine drift, and (f) Not-Clear, a burst-cluster with no clear pattern due to low signal-to-noise ratio or interference.
Further classification of ND burst-clusters has led to the identification of three sub-classes. ND-1 refers to single-component burst-clusters with a clear vertical pattern and a stable emission band, without frequency drifting ($\Delta\nu/\Delta t\approx\infty$). ND-2 and ND-m represent double and multiple component burst-clusters, respectively, which exhibit similar behavior but consist of two or more emission components that are arranged in parallel with approximately equal bandwidths ($\Delta\nu/\Delta t\approx 0$). The total number of ND burst-clusters is 720. Figure 7 illustrates examples of these sub-classes, where the first row showcases representative cases of ND-1, ND-2, and ND-m. The number of sub-classes is as follows: ND-1 (670 burst-clusters, 93.1%), ND-2 (49 burst-clusters, 6.8%), and ND-m (1 burst-cluster, 0.1%). Additionally, there are 9 consecutive time intervals and 42 intermittent time intervals for No Drifting burst-clusters.
3.4 burst-clusters with No Evidence for Drifting
3.5 burst-clusters with Not-Clear
3.6 Complex burst-clusters
These burst-clusters display fascinating features with complex structures in their waterfall plots, such as combinations of upward and downward drifts, or some sub-bursts within certain clusters exhibiting very intricate shapes. The diversity and intricacy of these burst-clusters make it worthwhile to review all the plots in extended material.
4 Statistical Relationship Between Drifting Rate and Other Physical Quantities
Downward drifting is a common characteristic of repeating FRBs observed across different telescopes and frequency bands. In FRB 20121102A, this phenomenon has been consistently reported in bursts spanning a wide frequency range (Hessels et al., 2019; Josephy et al., 2019; Platts et al., 2021). Similar drifting patterns were identified in FRB 20180814A, where sub-bursts exhibited progressive shifts to lower frequencies within the CHIME band (CHIME/FRB Collaboration et al., 2019b; Chawla et al., 2020). Observations of FRB 20180916B using simultaneous LOFAR (110–190 MHz) and Apertif (1220–1520 MHz) data confirmed that its characteristic sub-burst downward drifting behavior persists for a decade at radio frequency, with detections down to at least 120 MHz (Pastor-Marazuela et al., 2021), while CHIME observations revealed similar drifting structures in the 400–800 MHz band (Fonseca et al., 2020). While early studies of FRB 20190711 suggested downward drifting features (Day et al., 2020), subsequent analyses revealed ambiguities in burst classification. FRB 20201124A displayed complex time-frequency structures with consistent downward drifting across multiple observational campaigns (Hilmarsson et al., 2021; Zhou et al., 2022).
These studies demonstrate significant variations in drifting rates among different FRBs, potentially correlated with other burst parameters. Notably, Chamma et al. (2021) identified an inverse relationship between drifting rates and sub-burst durations, suggesting a unified mechanism governing these observed properties. Previous studies (Pleunis et al., 2021) reported drifting rates ranging from $-1$ to $-60$ MHz ms-1 in repeating FRB 20121102A and FRB 20180916B, and Zhou et al. (2022) reported drifting rates ranging from $-5$ to $-166$ MHz ms-1 in repeating FRB 20201124A. Our study of FRB 20240114A demonstrates that its drifting rates span an even broader range than previously reported, as shown in Figures 10, 11 and 12. Part of this can be attributed to the fact that our drifting rate measurements are based on burst-clusters, which may extend the range of drifting rates toward values closer to zero for the double- and multi-components burst-clusters. However, this approach does not affect the extremes of positive or negative infinity for all the burst-clusters.
4.1 Drifting Rate Analysis
Figure 8: Example of the drifting rate calculation for a single-component burst-cluster. The left panel displays the dynamic spectrum with the emission bandwidth evenly divided into three sub-bands. Gaussian fits are applied separately to each sub-band to determine the corresponding peak times and peak frequencies. The drifting rate is then computed using Orthogonal Distance Regression (ODR), a method that minimizes the sum of the squared orthogonal distances between the data points and the fitted model, thereby accounting for uncertainties in both time and frequency measurements. In this example, the measured drifting rate is $R_{d}=-156.90\pm 28.37\,\mathrm{MHz\,ms}^{-1}$.
Figure 9: Drifting rate calculation for a multi-components burst-cluster. For each sub-burst, an optimal time window is manually selected. Within this window, time integration and Gaussian fitting are performed to determine the peak frequency, and the drifting rate is subsequently derived by combining the peak times of the sub-bursts. In this example, the measured drifting rate is $R_{d}=-45.8895\pm 21.8814\,\mathrm{MHz\,ms}^{-1}$. A similar procedure is applied to double-component burst-clusters.
Following Zhou et al. (2022), we define the drifting rate as:
$R_{\rm d}=\frac{{\rm d}\nu}{{\rm d}t}.$ (3)
We categorize burst-clusters into single-, double-, and multi-components burst-clusters, applying different methodologies to determine their drifting rates.
For single-component burst-clusters, the emission bandwidth is evenly divided into three sub-bands. For each sub-band, Gaussian fitting is applied separately to determine both the peak time and peak frequency. The drifting rate is then calculated based on these values, as illustrated in Figure 8. For double- and multi-components burst-clusters, we manually select an appropriate time window for each sub-burst. Within the chosen window, time integration is performed, followed by a Gaussian fitting to extract the peak frequency. The drifting rate is then determined by combining the peak times of each pulse, as shown in Figure 9. It is important to note that the definition of drifting rate $R_{d}$ is more intuitive for the double- and multi-components burst-clusters, where the unit is MHz/ms, meaning that the larger the $R_{d}$, the more pronounced the drifting behavior. In contrast, for the single-component burst-clusters, when the drifting rate approaches infinity, it indicates that there is no drifting.
4.2 Statistical Relationships Between Peak Frequency, Drifting Rate, Bandwidth, Effective Width and Energy
Figure 10: This scatter matrix visualizes the relationships between the peak frequency, drifting rate, bandwidth, effective width, and energy for the single-peak dataset. The lower triangle contains pairwise scatter plots, and the diagonal displays the distribution of each variable. The different colors represent various morphological types (D1-H, D1-M, D1-L, D1-W for downward drifting and U1-H, U1-M, U1-L, U1-W for upward drifting), helping to differentiate the data samples.
Figure 11: Scatter plot of the drifting rate (MHz/ms) versus flux (Jy) for single-component burst-clusters exhibiting drifting behavior, with points color-coded by morphological classification (D1-H, D1-M, D1-L, D1-W for downward drifting and U1-H, U1-M, U1-L, U1-W for upward drifting). The plot shows that the flux of downward drifting events is slightly higher than that of upward drifting events.
In this section, we explore the statistical relationships between various physical parameters: (average) peak frequency, drifting rate, bandwidth, effective width, and isotropic energy. The format used is XY-Z, where X denotes the drifting type (D for downward drifting, U for upward drifting), Y represents the number of components (or sub-bursts) in the burst-cluster, with ”m” used if the number exceeds two, and Z indicates the observation band: L (low frequency), M (middle frequency), H (high frequency), W (wide frequency). For all the single-component burst-clusters exhibiting drifting behavior, these relationships are visualized using scatter matrices, which show pairwise relationships between variables along with the distribution of each individual variable along the diagonal as shown in 10, where only single-component burst-clusters data (Y = 1) are presented, with different values of X (D and U) and Z (H, M, L and W) represented by points of varying colors. Extending this analysis to double- and multi-components burst-clusters, the relationship between drifting rate and average peak frequency is visualized in Figure 12, with distinct color coding representing different combinations of the X and Z parameters. For the double- and multi-components burst-clusters datasets, the average peak frequency refers to the average of the individual peak frequencies for each sub-burst.
The U1 burst-clusters (single-component with positive drifting rate) exhibit a statistically significant linear correlation between peak frequency and minimum drifting rate, as shown in Figure 10. In contrast, neither the D1 clusters (single-component with negative drift, as seen in Figure 10) nor the Y=2/m composite clusters show such a dependence as demonstrated in Figure 12. The analysis presented here is based on the average DM for the day. Therefore, for single-component burst-clusters, the drifting rate may be affected by the accuracy of the DM correction. If the drifting rate for single-component burst-clusters could be fully corrected by an accurate DM, resulting in ND1, then an inaccurate DM correction would naturally lead to the drifting rate for high-frequency burst-clusters appearing higher compared to low-frequency burst-clusters, making the drifting less pronounced. This can be understood from equation (Zhang, 2023):
@@ 4.1 Drifting Rate Analysis
An inaccurate DM correction would result in a higher drifting rate at high frequencies. If the observed increase in drifting rate with peak frequency for U1 burst-cluster is attributed to this effect, a similar trend would be expected for D1. However, no such pattern is observed in Figure 10, suggesting that DM correction is just one of the factors that influence drifting behavior, and other physical mechanisms may also play a role. Therefore, the observed increase in drifting rate as the peak frequency increases for U1 burst-clusters remain notable.
As demonstrated in Figure 10, an inverse correlation between the drifting rate and sub-burst effective widths was robustly detected, in agreement with the findings of Chamma et al. (2021). Furthermore, for single-component burst-clusters, the D1 burst-clusters exhibit significantly larger bandwidth, effective width, and isotropic energy than U1, as shown in Figure 10. Considering that the isotropic energy is positively correlated with both the bandwidth and the effective width, we also plotted the relationship between the flux and the drifting rate. It can be seen that the flux of D1 events remains obviously higher than that of U1 events, as shown in Figure 11.
4.3 Analysis of consecutive and intermittent for downward and upward drifting

### v3

@@ III.2 Upward Drifting burst-clusters
In 2022RAA....22l4001Z, FRB 20201124A exhibited 263 downward drifting bursts (including single-, double-, and multi-component bursts) alongside 3 upward drifting bursts, all of which were double-component bursts. In contrast, our classification of FRB 20240114A shows that upward drifting burst-clusters account for 23.82% (233 out of 978) of the total burst-clusters exhibiting drifting behavior. However, the drifting mode of single-component burst-clusters is significantly affected by the choice of DM, while double- and multi-component burst-clusters are not subject to this influence. If we exclude single-component burst-clusters with upward drifting, the number of upward drifting double- and multi-component burst-clusters still reaches 91, accounting for 10.89% of the total sample of 836 burst-clusters exhibiting drifting behavior (excluding single-component burst-clusters with upward drifting). Furthermore, if we consider only upward drifting bursts, specifically those upward drifting burst-clusters consisting solely of consecutive time intervals, this number drastically decreases to just 9, as illustrated in Figure 6. This proportion is comparable to that observed in 2022RAA....22l4001Z. Upward drifting bursts, though rare, have theoretical grounding in magnetospheric trigger dynamics. Sparks triggered earlier may occur along field lines with a larger radius of curvature (2020ApJ...899..109W). Propagation effects such as interstellar scintillation can enhance the detection of upward drifting features through temporary intensity amplification (1990ARA&A..28..561R), although intrinsic emission mechanisms remain dominant.
III.3 No Drifting burst-clusters
@@
Figure 7: Representative examples of burst-cluster classifications. The first row shows No Drifting burst-clusters: (a) ND1, a single-component burst-cluster with a stable emission band and no frequency drift, (b) ND2, a double-component burst-cluster with two components showing no drift, and (c) NDm, a multi-component burst-cluster with parallel components. The second row illustrates burst-clusters with no evidence for drifting: (d) NE-H, a high-frequency burst-cluster with insufficient bandwidth to determine drift, (e) NE-L, a low-frequency burst-cluster with insufficient bandwidth to determine drift, and (f) Not-Clear, a burst-cluster with no clear pattern due to low signal-to-noise ratio or interference.
Further classification of ND burst-clusters has led to the identification of three sub-classes. ND1 burst refers to single-component burst-cluster with a clear vertical pattern and a stable emission band, without frequency drifting ($\Delta\nu/\Delta t\approx\infty$). However, as the drifting rate is affected by the value of DMs, a slightly different DM could cause the ND1 to transition to either downward or upward drifting with a high absolute drifting rate, and vise versa. One should be cautious regarding the statistics of the ND1 bursts. ND2 and NDm burst-clusters represent double- and multi-component burst-clusters, respectively, which exhibit similar behavior but consist of two or more emission components that are arranged in parallel with approximately equal bandwidths ($\Delta\nu/\Delta t\approx 0$). The total number of ND burst-clusters is 720. Figure 7 illustrates examples of these sub-classes, where the first row showcases representative cases of ND1, ND2, and NDm. The number of sub-classes is as follows: ND1 (670 burst-clusters, 93.1%), ND2 (49 burst-clusters, 6.8%), and NDm (1 burst-cluster, 0.1%). Additionally, there are 9 consecutive time intervals and 42 intermittent time intervals for No Drifting burst-clusters.
III.4 burst-clusters with No Evidence for Drifting
III.5 burst-clusters with Not-Clear
III.6 Complex burst-clusters
These burst-clusters display fascinating features with complex structures in their waterfall plots, such as combinations of upward and downward drifts, or some sub-bursts within certain bursts exhibiting very intricate shapes. The diversity and intricacy of these burst-clusters make it worthwhile to review all the plots in extended material.
IV Statistical relations Between Drifting Rate and Other Physical Quantities
Downward drifting is a common characteristic of repeating FRBs observed across different telescopes and frequency bands. In FRB 20121102A, this phenomenon has been consistently reported in bursts spanning a wide frequency range (Hessels_2019; 2019ApJ...882L..18J; 2021MNRAS.505.3041P). Similar drifting patterns were identified in FRB 20180814A, where sub-bursts exhibited progressive shifts to lower frequencies within the CHIME band (2019Natur.566..235C; 2020ApJ...896L..41C). Observations of FRB 20180916B using simultaneous LOFAR (110–190 MHz) and Apertif (1220–1520 MHz) data confirmed that its characteristic sub-burst downward drifting behavior persists for a decade of radio frequency, with detections down to at least 120 MHz (2021ApJ...911L...3P; 2021Natur.596..505P), while CHIME observations revealed similar drifting structures in the 400–800 MHz band (2020ApJ...891L...6F). While early studies of FRB 20190711 suggested downward drifting features (2020MNRAS.497.3335D), subsequent analyses revealed ambiguities in burst classification, mainly due to the burst’s complex temporal and spectral features, including multiple sub-bursts with varying polarization and scattering effects, which made it difficult to clearly classify it as a repeating or non-repeating burst. FRB 20201124A displayed complex time-frequency structures with consistent downward drifting across multiple observational campaigns (2021MNRAS.508.5354H; 2022RAA....22l4001Z).
These studies demonstrate significant variations in drifting rates among different FRBs, potentially correlated with other burst parameters. Notably, 2021MNRAS.507..246C identified an inverse relation between drifting rates and sub-burst durations, suggesting a unified mechanism governing these observed properties. Previous studies reported drifting rates of order 200-200 MHz ms-1 in repeating FRB 20121102A (2019ApJ...876L..23H), ranging from 1-1 to 60-60 MHz ms-1 in FRB 20180916B (2021ApJ...911L...3P), and from 5-5 to 166-166 MHz ms-1 in FRB 20201124A (2022RAA....22l4001Z). Our study of FRB 20240114A demonstrates that its drifting rates span an even broader range than previously reported. Part of this can be attributed to the fact that our drifting rate measurements are based on burst-clusters, which may extend the range of drifting rates toward values closer to zero for the double- and multi-component burst-clusters, as discussed in Section IV.1. However, this approach does not affect the extremes of positive or negative infinity for all the burst-clusters.
IV.1 Drifting Rate Analysis
Figure 8: Example of the drifting rate calculation for a single-component burst-cluster. The left panel displays the dynamic spectrum with the emission bandwidth evenly divided into three sub-bands. Gaussian fits are applied separately to each sub-band to determine the corresponding peak times and peak frequencies. The drifting rate is then computed using ODR, a method that minimizes the sum of the squared orthogonal distances between the data points and the fitted model, thereby accounting for uncertainties in both time and frequency measurements. In this example, the measured drifting rate is $R_{d}=-156.90\pm 28.37\,\mathrm{MHz\,ms}^{-1}$.
Figure 9: Drifting rate calculation for a multi-component burst-cluster. For each sub-burst, an optimal time window is manually selected. Within this window, time integration and Gaussian fitting are performed to determine the peak frequency, and the drifting rate is subsequently derived by combining the peak times of the sub-bursts. In this example, the measured drifting rate is $R_{d}=-45.89\pm 21.88\,\mathrm{MHz\,ms}^{-1}$. A similar procedure is applied to double-component burst-clusters.
Following 2022RAA....22l4001Z, we define the drifting rate as:
$R_{\rm d}=\frac{{\rm d}\nu}{{\rm d}t}.$ (3)
We categorize burst-clusters into single-, double-, and multi-component burst-clusters, applying different methodologies to determine their drifting rates. The drifting of single-component bursts can be considered as intra-burst drifting, while the drifting of both the double- and multi-component burst-clusters can be considered as inter-burst drifting.
For single-component burst-clusters, the emission bandwidth is evenly divided into three sub-bands as a sampling strategy to obtain multiple independent frequency–time centroids. Within each sub-band, Gaussian fitting is applied to determine both the peak time and peak frequency. The drifting rate is then derived from these values using Orthogonal Distance Regression (ODR), as illustrated in Figure 8. For double- and multi-component burst-clusters, we manually select an appropriate time window for each sub-burst. Within the chosen window, time integration is performed, followed by a Gaussian fitting to extract the peak frequency. The drifting rate is then determined by combining the peak times of each sub-burst, as shown in Figure 9. It is important to note that the drifting rate $R_{d}$ is most intuitive for double- and multi-component burst-clusters, where the unit is MHz ms-1 and a larger $|R_{\rm d}|$ corresponds to a more pronounced drifting behavior. In contrast, for the single-component burst-clusters, when the drifting rate approaches infinity, it indicates that there is no drifting.
IV.2 Statistical relations Between Peak Frequency, Drifting Rate, Bandwidth, Effective Width and Energy
Figure 10: This scatter matrix visualizes the relations between the peak frequency, drifting rate, bandwidth, effective width, and energy for the single-peak dataset. The lower triangle contains pairwise scatter plots, and the diagonal displays the distribution of each variable. The different colors represent various morphological types (D1-H, D1-M, D1-L, D1-W for downward drifting and U1-H, U1-M, U1-L, U1-W for upward drifting), helping to differentiate the data samples.
Figure 12: Scatter plots of the drifting rate with frequency peak (left panel) and flux (right panel) for single-component burst-clusters exhibiting measurable frequency drifts. Each point represents an individual burst-cluster, color-coded by morphological subtype (D1-H, D1-M, D1-L, D1-W for downward drifting and U1-H, U1-M, U1-L for upward drifting). ODR fits for D1 and U1 are plotted in gray (left panel). In the right panel, horizontal gray lines mark the mean fluxes of the two classes, $\langle F_{\mathrm{D1}}\rangle=0.35\pm 0.36~\mathrm{Jy}$ and $\langle F_{\mathrm{U1}}\rangle=0.16\pm 0.17~\mathrm{Jy}$, indicating that D1 burst-clusters are on average about 2.2 times brighter. Error bars are plotted only for bursts with frequency peak and drifting rate uncertainties below 500 MHz and 400 MHz ms-1, respectively.
In this section, we explore the statistical relations between various physical parameters: (average) peak frequency, drifting rate, bandwidth, effective width, and isotropic energy. The format used is XY-Z, where X denotes the drifting type (D for downward drifting, U for upward drifting), Y represents the number of components (or sub-bursts) in the burst-cluster, with ”m” used if the number exceeds two, and Z indicates the observation band: L (low frequency), M (middle frequency), H (high frequency), W (wide frequency). For all the single-component burst-clusters exhibiting drifting behavior, these relations are visualized using scatter matrices, which show pairwise relations between variables along with the distribution of each individual variable along the diagonal as shown in Figure 10, where only single-component burst-clusters data (Y = 1) are presented, with different values of X (D and U) and Z (H, M, L and W) represented by points of varying colors. Extending this analysis to double- and multi-component burst-clusters, the relation between drifting rate and average peak frequency is visualized in Figure 11, with distinct color coding representing different combinations of the X and Z parameters. For the double- and multi-component burst-clusters datasets, the average peak frequency refers to the average of the individual peak frequencies for each sub-burst.
The U1 and D1 burst-clusters (single-component with drifting behavior) exhibit a clear correlation between peak frequency and drifting rate, as shown in Figures 10 and 12, while the Y=2/m burst-clusters also lack such a correlation, as shown in Figure 11.
We make the simplifying assumption that the measured drifting rate is entirely due to the deviation between the adopted average DM and the true per-burst DM, as described by the cold-plasma dispersion relation (2023RvMP...95c5005Z):
$\Delta t=t(\nu_{1})-t(\nu_{2})=\frac{e^{2}}{2\pi m_{e}c}\!\left(\frac{1}{\nu_{1}^{2}}-\frac{1}{\nu_{2}^{2}}\right)\!\mathrm{DM}\approx 4.15\,\mathrm{ms}\!\left(\frac{1}{\nu_{1,\mathrm{GHz}}^{2}}-\frac{1}{\nu_{2,\mathrm{GHz}}^{2}}\right)\frac{\mathrm{DM}}{\mathrm{pc}\,\mathrm{cm}^{-3}}.$ (4)
where $\Delta t$ represents the time delay between two frequencies $\nu_{1}$ and $\nu_{2}$, with $e$ being the electron charge, $m_{e}$ the electron mass, and $c$ the speed of light. The term $\frac{1}{\nu_{1}^{2}}-\frac{1}{\nu_{2}^{2}}$ accounts for the difference in arrival times at different frequencies, while DM quantifies the integrated free electron density along the line of sight.
Taking the derivative shows that, under the assumption that DM deviations are the only source of the observed slope, the measured drifting rate should increase in magnitude with frequency. In our analysis the daily average DM is used rather than the exact DM of each burst (assuming such measurements could be accurate). In this case, both upward and downward drifting burst-clusters are expected to follow the same frequency-dependent scaling in $|R_{d}|$. However, our results reveal that the two classes exhibit significantly different slopes: U1 burst-clusters show a strong positive correlation (ODR slope $b=1.20\pm 0.07~\mathrm{ms}$; Pearson $r=0.47$, $p=3.5\times 10^{-9}$; Spearman $\rho=0.54$, $p=2.7\times 10^{-12}$), while D1 burst-clusters show a weaker dependence (ODR slope $b=-2.86\pm 0.24~\mathrm{ms}$; Pearson r=0.21r=-0.21, $p=9.0\times 10^{-5}$; Spearman $\rho=-0.23$, $p=7.6\times 10^{-6}$), as shown in the left panel of Figure 12. The substantial difference in slope magnitude between U1 and D1 indicates that DM deviations alone cannot account for the observed trend, indicating that additional, yet unidentified physical mechanisms may contribute to the frequency dependence of the drifting rate. Furthermore, for single-component burst-clusters, the D1 burst-clusters exhibit significantly larger bandwidth, effective width, and isotropic energy than U1, as shown in Figure 10. It is interesting to see that the energy for downward drifting bursts is statistically higher than the energy for the upward drifting ones (lower-left panel). Considering that the isotropic energy of detected bursts can be calculated from the emission bandwidth and effective width, we also plotted the relation between the flux and the drifting rate. The average flux of D1 burst-clusters ($\langle F_{\mathrm{D1}}\rangle=0.35\pm 0.36~\mathrm{Jy}$) is about 2.19 times higher than that of U1 burst-clusters ($\langle F_{\mathrm{U1}}\rangle=0.16\pm 0.17~\mathrm{Jy}$), as shown in the right panel of Figure 12.
Another physically meaningful statistical parameter is $\Delta\nu/\nu_{0}$, which characterizes the degree of spectral confinement of the emission. Here, $\Delta\nu$ denotes the full width at half-maximum (FWHM) of the emission spectrum and $\nu_{0}$ the central frequency, following the definition in 2024ApJ...974..160K. In our analysis, the bandwidths are adopted from the catalog of 2025arXiv250714707Z, where the frequency boundaries were determined using a CDF-based edge-finding method, with Gaussian extrapolation applied to band-truncated bursts and manual selection for a few complex or low-S/N cases. These dataset bandwidths are primarily defined as $2\sigma$ (or CDF-based equivalent) bandwidths, and the Gaussian and CDF results are found to be generally consistent within the measurement uncertainties. For Gaussian profiles, the $2\sigma$ width can be approximately converted to the equivalent FWHM by a factor of $\sqrt{2\ln 2}\approx 1.18$.
For the bursts detected on March 12, 2024, the histogram of $\Delta\nu/\nu_{0}$ is well fitted by a Gaussian, with the full dataset yielding $\mu=0.223$ and $\sigma=0.093$, and the frequency-limited subset ($1.05~\mathrm{GHz}<\nu_{0}\pm\Delta\nu/2<1.45~\mathrm{GHz}$) fitted by a Gaussian with $\mu=0.187$ and $\sigma=0.053$, as shown in Figure 13. Since many bursts exhibit spectral components extending beyond the 1–1.5,GHz FAST band, focusing on those fully enclosed within the band ensures that the measured bandwidths are not limited by the instrumental frequency coverage. After applying the FWHM conversion, the typical fractional bandwidths become $\Delta\nu/\nu_{0}\approx 0.22$-$0.27$, still well below the theoretical lower limit of about $\Delta\nu/\nu_{0}\sim 0.58$, which is expected for emission outside the magnetosphere due to high-latitude broadening (2024ApJ...974..160K). This lower bound arises because, for an FRB source outside the magnetosphere whose angular size exceeds the Doppler beaming angle, high-latitude emission causes spectral broadening that sets a minimum fractional bandwidth. Such high-latitude effects are therefore difficult to reconcile with the markedly confined spectra observed here. This strong spectral confinement, therefore, favors a magnetospheric emission origin; however, additional propagation effects, such as plasma lensing or scintillation, may still influence the observed spectral morphology. Similar narrow-band behavior has also been noted in previous studies of repeating FRBs, including FRB 20220912A, FRB 20230607A, and FRB 20240114A, whose fractional bandwidth distributions typically peak around $\Delta\nu/\nu_{0}\sim 0.1$-$0.2$ (2025ApJ...988...41Z; 2023ApJ...955..142Z; 2024ApJ...977..177K).
Figure 13: Distribution of $\Delta\nu/\nu_{0}$ for FRB 20240114A bursts detected on March 12, 2024. The blue and orange histograms represent the full dataset and the frequency-limited subset ($1.05~\mathrm{GHz}<\nu_{0}\pm\Delta\nu/2<1.45~\mathrm{GHz}$), respectively. The solid green and red curves show the corresponding Gaussian fits, yielding $(\mu,\,\sigma)=(0.223,\,0.093)$ and (0.187, 0.053)(0.187,\,0.053). After converting the dataset $2\sigma$ bandwidths to the equivalent FWHM ($\mathrm{FWHM}/(2\sigma)=\sqrt{2\ln 2}\approx 1.18$), the typical fractional bandwidths become $\Delta\nu/\nu_{0}\approx 0.22$$0.27$, still well below the theoretical lower limit of about $\Delta\nu/\nu_{0}\sim 0.58$ expected for emission outside the magnetosphere.
IV.3 Analysis of consecutive and intermittent for downward and upward drifting